The perimeter of a rectangular garden is 24 ft. The length is 2 ft more than the width. Find the length and the width of the garden
Answer by math problem solving(30)

Let the width be x.
Then the length is x + 2.
The perimeter of the rectangle in terms of x is:
x + x + (x + 2) + (x + 2) = 4x + 4
As the given measure of the perimeter is 24
we have
4x + 4 = 24
Solving for x, we have
4x = 20
x = 5
So the width is 5 ft, and the length is x + 2= 5 + 2 = 7 ft.